OpenStudy (somegirl1999):

Tamara needs to buy motor oil to fill the 3 empty cylindrical barrels at her oil service center. Each barrel is 7 ft deep and has a radius of 4 ft. What is the volume of oil needed? Use CE_pi_large = 3.14. (1 point) 1,055.04 ft3 4,220.16 ft3 602.88 ft3 1,406.72 ft3

OpenStudy (anonymous):

what is the formula for Volume of a cylinder?

OpenStudy (somegirl1999):

um to x's the base and the height ??

OpenStudy (somegirl1999):

im sorry im dumb when it comes to math lol

OpenStudy (anonymous):

sry im back :)

OpenStudy (anonymous):

You didnt understand when i explained it last time? you still have that last post?

OpenStudy (somegirl1999):

lol i couldnt put it in the calculator xD.

OpenStudy (anonymous):

haha ok ill try and explain it better this time :)

OpenStudy (somegirl1999):

ok lol thanks

OpenStudy (somegirl1999):

i need 1 on 1 on math my dad is the math genius in our family

OpenStudy (anonymous):

the area of a cylinder is the area of a circle multiplied by the height. therefore \[h(\pi ^{2})\]

OpenStudy (anonymous):

oh ok well you can always come here and ask me :D

OpenStudy (anonymous):

that is if i can answer them xD

OpenStudy (somegirl1999):

lol ok thanks

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i wrote pellet xD

OpenStudy (somegirl1999):

lolol

OpenStudy (anonymous):

\[h(\pi r ^{2})\]

OpenStudy (anonymous):

lol they cencored my colourful language :P

OpenStudy (somegirl1999):

lolol

OpenStudy (anonymous):

h=7 \[\pi = 3.14\] r=4

OpenStudy (anonymous):

if you want u can do your woking here and ill correct it but if not then dont forget the multiply your answer by 3 cause there r 3 barrells

OpenStudy (somegirl1999):

ok

OpenStudy (anonymous):

cool :D lol ill try not to tell you more pelletz

OpenStudy (somegirl1999):

Lmao

OpenStudy (anonymous):

haha

OpenStudy (somegirl1999):

brb need a drink haha

OpenStudy (anonymous):

lol np

OpenStudy (somegirl1999):

ok im back :P

OpenStudy (anonymous):

wb :p

OpenStudy (somegirl1999):

alright so on the question i do 7x3.14x4 and thanks :P

OpenStudy (anonymous):

4^2

OpenStudy (anonymous):

so 7x3.14x4^2

OpenStudy (somegirl1999):

ok thank you

OpenStudy (anonymous):

then multiply ur answer by 3 cause there r 3 barrels

OpenStudy (somegirl1999):

omg thank you so much :)

OpenStudy (anonymous):

ur welcome :)